OpenStudy (anonymous):

Find the chance that the 5 people have 5 different birthdays. Given that: a) There are no twins no leap years b) One year has 365 days

OpenStudy (anonymous):

HI it would be b

OpenStudy (anonymous):

Pardon

OpenStudy (anonymous):

Is that the full question?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

b

OpenStudy (anonymous):

The chance a single person will have their birthday on any day of the year would be (1/365). I remember learning how to calculate the chances of multiple people getting different value, but I forgot how to do it. Let me do some research.

OpenStudy (anonymous):

Let me tell you all that a and b are not the choices to the answer of the question but the condition to solve the problem

OpenStudy (anonymous):

ok take your time

OpenStudy (anonymous):

Mmk, it's "probability when order matters". So it would be something along the lines of: \[\frac{ \left( 5 \times 4 \times 3 \times 2 \times 1 \right) }{ 365 }\]

OpenStudy (anonymous):

Is it on edx.org? Have seen SOOO MANY people asking the same questions over and over again and sadly, all of them want nothing but the answers.

OpenStudy (anonymous):

Don't know what you talking about drawar but I am stuck with this question of the assignment

OpenStudy (agent0smith):

The first person can have any birthday out of the 365 days, the second person only has 364 possible days (since they can't have the same day as the first person), then 363 possible days for the next person and so on \[\large \frac{ 365 \times 364 \times 363\times 362 \times 361 }{ 365 \times 365 \times365 \times365 \times365 }\]

OpenStudy (anonymous):

365×364×363×362×361/365×365×365×365×365 =364x363x362x361/365^4 =0.9728644263 =97.28644263%

Parth (parthkohli):

The answer is b

OpenStudy (shubhamsrg):

C(365,5)/(365^5)

OpenStudy (shubhamsrg):

that'll be multiplied by 5! so final asnwer would be what @agent0smith says also, b) is right ^_^ :D